a gas mixture of 3lits of propane and butane on complete combustion at 298kproduce 10 lits of carbondioxide. find out composition of gas mixture is?

Grade:Select Grade

1 Answers

Sunil Kumar FP
9 years ago
let the amount of propane be x lit and the amount of butane be 3-x lt.
C3H8 +O2 ---->3CO2 +4H2O
C4H10 +13/2O2 ------->4CO2 +5H2O
x lt will give 3x lt of CO2 and 3-x lt will give 4(3-x)lt of CO2
Total amount of CO2 produced=3x+4(3-x)lt
=(12-x) lt
we have total amount of CO2 produced =12-x lt
12-x=10 lt
x=2lt
Amount of propane= 2lt and amount of butane=1lt
%of propane=2/3 *100 =66.66%
% of butane =1/3 *100 =33.33%